# Coarse-graining particulate two-phase flow

### v7

@@ 2.2.3 Reynolds rule, uniqueness, and coarse-graining scale
Otherwise, like for filtering procedures used in Large Eddy Simulation [26], the resulting coarse-grained fields are not unique as they depend on the details of the averaging procedure. Excluding the trivial (and generally unphysical) cases of the identity operator and operators that average over all times and/or space, Eq. (12) essentially states that the averaging should include the entire ensemble of microscopic states belonging to a given macroscopic state, while not intermingling different macroscopic states. There are only three kind of standard averaging procedures that identically satisfy this rule: the ensemble average, the infinite time average in steady systems, and infinite averages over a given spatial direction if a system is statistically homogeneous in that direction. Assuming that the ergodic hypothesis is true, all these three averages are equivalent, and averaged quantities are therefore unique. However, in practice, ensemble averaging is inconvenient and computationally costly, and systems are generally statistically heterogeneous in space and time. For this reason, one often resorts to sphere-symmetrical spatial averaging over a characteristic coarse-graining scale $\zeta$ [6, 15]. This scale should be sufficiently large to capture sufficiently many microscopic states, but not too large to substantially intermingle macroscopic states, which essentially necessitates scale separation between the important microscopic and macroscopic scales [6]. In this regard, the important microscopic scales are often better represented by the typical separation distance between particles than by $R$ [27], implying that, in systems with high solid volume fraction $\beta_{\mathrm{s}}$, such as dense granular flows, this condition might be satisfied even when scale separation between $R$ and $L$ is lacking. Then, there is often a goldilock zone, a range of $\zeta$ where sphere-symmetrically spatially averaged fields are almost unique [15], which is precisely the zone where such averages approximately satisfy the Reynolds rule.
@@ 2.3 Derivation of macroscopic balance equations
The derivations presented in this section follow a simple strategy. First, a balance equation is formulated at a microscopic, distributional level, involving generalized functions such as $X^{p}$ and $\delta[\bm{x}-\bm{x}^{p}]$. Afterward, a generic averaging operator $\langle\cdot\rangle$ that satisfies the rules in Eqs. (7) and (12) is applied to both sides of the equation. The commutation rule then permits exchanging the order of differentiation and averaging, which leads to a macroscopic balance equation via exploiting the phase average definitions in Eqs. (9)-(11) and, if required, via carrying out a Reynolds decomposition.
@@ 2.3.3 Macroscopic momentum balance equations
Summing, averaging: $\langle\sum_{p}\cdot\rangle$ Averaging rulesPhase average definitionsReynolds decomposition Express $X^{p}\bm{f}^{p}_{\mathrm{c}}$ $\bm{\nabla}\cdot(\dots), as$ shown in Eqs. (34-35)
@@ 2.3.4 Note on Reynolds stresses and fluctuation velocities
When numerical implementing the coarse-graining method, one should use the top-line expressions in Eqs. (24), (38), and (42) to calculate the Reynolds stresses rather than the usual respective bottom-line expressions based on the fluctuation velocities, since the latter require the Reynolds rule to be valid, which is almost never identically satisfied in practice. This should automatically resolve previously tackled issues regarding a grid-independent definition of the fluctuation velocity [16]. In fact, in the fluid mechanics community, the analog of the top line of Eq. (24) has been used for several decades to calculate the pure-fluid Reynolds stress tensor in numerical simulations [2].
@@ 3.1.1 IBM after Uhlmann (2005)
In order to adapt the formalism developed in §2 to Eq. (44), one first needs to circumvent the mathematical problem that the generalized-function products $X^{p}[\bm{x}]\delta[\bm{x}-\bm{y}]$ and $X_{\mathrm{f}}[\bm{x}]\delta[\bm{x}-\bm{y}]$ are not well-defined for $\bm{y}\in\mathbb{S}^{p}$. To this end, we introduce an infinitesimally extended particle $p$, occupying a domain $\mathbb{V}^{p}_{+}$ that is constructed such that the surface $\mathbb{S}^{p}$ of $\mathbb{V}^{p}$ lies strictly in the interior of $\mathbb{V}^{p}_{+}$. For example, for a spherical particle $p$ of radius $R_{p}$, integration over the volume $\mathbb{V}^{p}_{+}$ would then be equivalent to integration over a spherical domain of radius $(1+\epsilon)R_{p}$ and performing the limit $\epsilon\rightarrow 0^{+}$ afterward. Likewise, we define the infinitesimally contracted domain $\mathbb{V}_{\mathrm{f}-}$ such that $\mathbb{S}^{p}$ lies strictly outside of $\mathbb{V}_{\mathrm{f}-}$. The corresponding generalized-function products therefore satisfy $X^{p}_{+}[\bm{x}]\delta[\bm{x}-\bm{y}]=\delta[\bm{x}-\bm{y}]$ and $X_{\mathrm{f}-}[\bm{x}]\delta[\bm{x}-\bm{y}]=0$, respectively, for $\bm{y}\in\mathbb{S}^{p}$. These relations imply that $\bm{f}_{\mathrm{IBM}}$ does not appear as an additional body force term in the macroscopic fluid phase momentum balance and that the hydrodynamic force acting on a particle $p$, using Eq. (44) and Reynolds transport theorem, can be calculated as [20]
@@
where $\mathbb{S}^{p}_{+}=\partial\mathbb{V}^{p}_{+}$ is the extended surface domain of a particle $p$ and $\bm{n}^{p}_{+}$ its outward-directed unit normal vector. The bottom line of Eq. (45) is precisely how $\bm{F}^{p}_{\mathrm{h}}$ is calculated in Uhlmann’s IBM [17, 18, 19, 20, 21], considering that these studies modeled the surface $\mathbb{S}^{p}$ as a very thin volume layer, the so-called Lagrangian layer [20], extending minimally both into the interior and exterior of the particle $p$. This layer was then discretized into small grid cells, so that in total:
@@
where $\bm{f}^{p}_{l}\Delta V^{p}_{l}$ is the force applied on the grid cell volume $\Delta V^{p}_{l}$ corresponding to a given Lagrange point (index $l$) on the surface of the particle $p$. In Eq. (46), and also in the equations of Appendix C, the symbol $\simeq$’ shall indicate approximations associated with DNS- or IBM-related discretizations.
@@
Here $\mathcal{G}_{\Delta}[\bm{x}]$ is a smoothing function that satisfies $\int_{\mathbb{R}^{3}}\mathcal{G}_{\Delta}[\bm{x}]\mathrm{d}x^{3}=1$ and exhibits a compact support on the order of the DNS grid length $\Delta$ (Appendix B). The key problem now is that $\mathcal{G}_{\Delta}$ spreads the particle surface force density $\bm{\tilde{f}}^{p}_{\mathrm{IBM}}$ onto the stationary Eulerian DNS grid, independently of $\Delta$ about equally to the interior and exterior of $\mathbb{V}^{p}_{+}$, with the result that $|\int_{\mathbb{V}^{p}_{+}}\overline{\bm{f}}_{\mathrm{IBM}}\mathrm{d}V|$ substantially underestimates $|\int_{\mathbb{V}^{p}_{+}}\bm{f}_{\mathrm{IBM}}\mathrm{d}V|$, leading to potentially substantial differences between $\int_{\mathbb{S}^{p}_{+}}\bm{n}^{p}_{+}\cdot\bm{\tilde{\sigma}}\mathrm{d}S$ and $\bm{F}^{p}_{\mathrm{h}}$. This issue remains present even in the limit $\Delta\rightarrow 0^{+}$, implying that it cannot be dismissed as a discretization error. It suggests that Eq. (47) may not be well approximated when the unknown $\bm{\sigma}$ on its right-hand side is approximated by $\bm{\tilde{\sigma}}$. In addition, in order to evaluate Eq. (47), one would then need to interpolate the values of $\bm{\tilde{\sigma}}$ back from the DNS grid to the particle surfaces, creating another potential source of inaccuracy.
@@
To solve both problems, we exploit the fact that the fluid fields are well-defined in the entire fluid-particle domain $\mathbb{V}_{\infty}$. Using the relations derived in Appendix A and $\bm{\nabla}\cdot\bm{\sigma}=\bm{\nabla}\cdot\bm{\tilde{\sigma}}+\overline{\bm{f}}_{\mathrm{IBM}}-\bm{f}_{\mathrm{IBM}}$, this permits calculating
@@
$\begin{split}&\bm{\nabla}\cdot S^{p}\underline{\bm{r}^{p}\bm{n}^{p}_{+}\cdot\bm{\sigma}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{S}^{p}_{+}}\\ =&\;\bm{\nabla}\cdot\left(X^{p}_{+}\bm{\sigma}+V^{p}\underline{\bm{r}^{p}(\bm{\nabla}\cdot\bm{\sigma})\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{V}^{p}_{+}}\right)\\ =&\;\bm{\nabla}\cdot\left(X^{p}\bm{\sigma}+V^{p}\underline{\bm{r}^{p}\left(\bm{\nabla}\cdot\bm{\tilde{\sigma}}+\overline{\bm{f}}_{\mathrm{IBM}}\right)\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{V}^{p}}\right.\\ &\quad\quad\quad\quad\;\;\,-\left.X^{p}_{+}\sum\nolimits_{q}S^{q}\underline{\bm{r}^{p}\bm{\tilde{f}}^{q}_{\mathrm{IBM}}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{S}^{q}}\right),\end{split}$ (50)
@@
where the term in the bottom-most line accounts for the fact that a portion of the surface $\mathbb{S}^{q}$ of a particle $q$ can be within the extended domain $\mathbb{V}^{p}_{+}$ of a particle $p$ not only if $q=p$ but also if $q\neq p$, since contacting particles are allowed to slightly overlap. Furthermore, since $\mathcal{G}_{\Delta}$ integrates to unity, the difference $\bm{\nabla}\cdot\bm{\sigma}-\bm{\nabla}\cdot\bm{\tilde{\sigma}}$ can be directly calculated from the surface force density $\bm{\tilde{f}}^{p}_{\mathrm{IBM}}$ as
@@
$\bm{\sigma^{\mathrm{f}}}=\langle\bm{\tilde{\sigma}}\rangle+\bm{\sigma^{\mathrm{f}}_{\mathrm{Re}}}+\bm{\sigma^{\mathrm{f}}_{\mathrm{\tilde{s}}}}+\bm{\sigma^{\mathrm{f}}_{\Delta}},$ (52)
$\begin{split}&\bm{\sigma^{\mathrm{f}}_{\mathrm{\tilde{s}}}}\equiv\left\langle\sum\nolimits_{p}\left(V^{p}\underline{\bm{r}^{p}\left(\bm{\nabla}\cdot\bm{\tilde{\sigma}}+\overline{\bm{f}}_{\mathrm{IBM}}\right)\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{V}^{p}}-\right.\right.\\ &\left.\left.X^{p}_{+}\sum\nolimits_{q}S^{q}\underline{\bm{r}^{p}\bm{\tilde{f}}^{q}_{\mathrm{IBM}}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{S}^{q}}-\underline{\bm{r}^{p}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{V}^{p}}\bm{F}^{p}_{\mathrm{h}}\right)\right\rangle,\end{split}$ (53)
@@
$\bm{\sigma^{\mathrm{m}}}=\langle\bm{\tilde{\sigma}}\rangle+\bm{\sigma^{\mathrm{f}}_{\mathrm{\tilde{s}}}}+\bm{\sigma^{\mathrm{f}}_{\Delta}}+\bm{\sigma^{\mathrm{s}}_{\mathrm{c}}}+\bm{\sigma^{\mathrm{m}}_{\mathrm{Re}}}.$ (55)

### v8

@@ 2.2.3 Reynolds rule, uniqueness, and coarse-graining scale
Otherwise, like for filtering procedures used in Large Eddy Simulation [26], the resulting coarse-grained fields are not unique as they depend on the details of the averaging procedure. Excluding the trivial (and generally unphysical) cases of the identity operator and operators that average over all times and/or space, Eq. (12) essentially states that the averaging should include the entire ensemble of microscopic states contributing to a macroscopic state at a given $\bm{x}$ and $t$, while not intermingling with different macroscopic states at surrounding $\bm{x}$ and $t$. There are only three kind of standard averaging procedures that identically satisfy this rule: the ensemble average, the infinite time average in steady systems, and infinite averages over a given spatial direction if a system is statistically homogeneous in that direction. Assuming that the ergodic hypothesis is true, all these three averages are equivalent, and averaged quantities are therefore unique. However, in practice, ensemble averaging is inconvenient and computationally costly, and systems are generally statistically heterogeneous in space and time. For this reason, one often resorts to sphere-symmetrical spatial averaging over a characteristic coarse-graining scale $\zeta$ [6, 15]. This scale should be sufficiently large to capture sufficiently many microscopic states, but not too large to substantially intermingle macroscopic states, which essentially necessitates scale separation between the important microscopic and macroscopic scales [6]. In this regard, the important microscopic scales are often better represented by the typical separation distance between particles than by $R$ [27], implying that, in systems with high solid volume fraction $\beta_{\mathrm{s}}$, such as dense granular flows, this condition might be satisfied even when scale separation between $R$ and $L$ is lacking. Then, there is often a goldilock zone, a range of $\zeta$ where sphere-symmetrically spatially averaged fields are almost unique [15], which is precisely the zone where such averages approximately satisfy the Reynolds rule.
@@ 2.3 Derivation of macroscopic balance equations
The derivation presented in this section follows a simple strategy. First, a balance equation is formulated at a microscopic, distributional level, involving generalized functions such as $X^{p}$ and $\delta[\bm{x}-\bm{x}^{p}]$. Afterward, a generic averaging operator $\langle\cdot\rangle$ that satisfies the rules in Eqs. (7) and (12) is applied to both sides of the equation. The commutation rule then permits exchanging the order of differentiation and averaging, which leads to a macroscopic balance equation via exploiting the phase average definitions in Eqs. (9)-(11) and, if required, via carrying out a Reynolds decomposition.
@@ 2.3.3 Macroscopic momentum balance equations
Summing, averaging: $\langle\sum_{p}\cdot\rangle$ Averaging rulesPhase average definitionsReynolds decomposition Express $X^{p}\bm{f}^{p}_{\mathrm{c}}$ $\bm{\nabla}\cdot(\dots), as$ shown in Eqs. (34)-(35)
@@ 2.3.4 Note on Reynolds stresses and fluctuation velocities
When numerical implementing the coarse-graining method, one should use the top-line expressions in Eqs. (24), (38), and (42) to calculate the Reynolds stresses rather than the usual respective bottom-line expressions based on the fluctuation velocities, since the latter require the Reynolds rule to be valid, which is almost never identically satisfied in practice. In fact, in the fluid mechanics community, the analog of the top line of Eq. (24) has been used for several decades to calculate the pure-fluid Reynolds stress tensor in numerical simulations [2].
Artoni and Richard [16] introduced an alternative definition of the fluctuation velocity for granular flows that eliminates the grid dependence of the solid phase Reynolds stress (based on the traditional microscopic volume fraction of a particle $p$, $\phi^{p}_{\mathrm{mic}}=V^{p}\delta[\bm{x}-\bm{x}^{p}]$) for systems with slowly-varying shear rate. However, since our coarse-graining method requires the Reynolds rule to be at least approximately satisfied, which implies almost no grid dependence of the coarse-grained fields (§2.2.3), there is no good justification for deviating from using Eq. (38) to calculate $\bm{\sigma^{\mathrm{s}}_{\mathrm{Re}}}$.
@@ 3.1.1 IBM after Uhlmann (2005)
In order to adapt the formalism developed in §2 to Eq. (44), one first needs to circumvent the mathematical problem that the generalized-function products $X^{p}[\bm{x}]\delta[\bm{x}-\bm{y}]$ and $X_{\mathrm{f}}[\bm{x}]\delta[\bm{x}-\bm{y}]$ are not well-defined for $\bm{y}\in\mathbb{S}^{p}$. To this end, we introduce an infinitesimally extended particle $p$, occupying a domain $\mathbb{V}^{p}_{+}$ that is constructed such that the surface $\mathbb{S}^{p}$ of $\mathbb{V}^{p}$ lies strictly in the interior of $\mathbb{V}^{p}_{+}$. For example, for a spherical particle $p$ of radius $R_{p}$, integration over the volume $\mathbb{V}^{p}_{+}$ would then be equivalent to integration over a spherical domain of radius $(1+\epsilon)R_{p}$ and performing the limit $\epsilon\to 0^{+}$ afterward. Likewise, we define the infinitesimally contracted domain $\mathbb{V}_{\mathrm{f}-}$ such that $\mathbb{S}^{p}$ lies strictly outside of $\mathbb{V}_{\mathrm{f}-}$. The corresponding generalized-function products therefore satisfy $X^{p}_{+}[\bm{x}]\delta[\bm{x}-\bm{y}]=\delta[\bm{x}-\bm{y}]$ and $X_{\mathrm{f}-}[\bm{x}]\delta[\bm{x}-\bm{y}]=0$, respectively, for $\bm{y}\in\mathbb{S}^{p}$. These relations imply that $\bm{f}_{\mathrm{IBM}}$ does not appear as an additional body force term in the macroscopic fluid phase momentum balance and that the hydrodynamic force acting on a particle $p$, using Eq. (44) and Reynolds transport theorem, can be calculated as [^1] The equality in the second line of Eq. (45) assumes $\int_{\mathbb{V}^{p}_{+}}\sum_{q}\int_{\mathbb{S}^{q}}\delta[\bm{x}-\bm{y}]\bm{\tilde{f}}^{q}_{\mathrm{IBM}}[\bm{y}]\mathrm{d}S_{y}\mathrm{d}V_{x}=\int_{\mathbb{S}^{p}}\bm{\tilde{f}}^{p}_{\mathrm{IBM}}\mathrm{d}S$, which neglects the fact that particles are allowed to very slightly overlap. The precise calculation would give nonzero contributions not only for $q=p$ but also for $q\neq p$ if a portion of the surface $\mathbb{S}^{q}$ of a particle $q$ is within the extended domain $\mathbb{V}^{p}_{+}$ of a particle $p$.
@@
where $\mathbb{S}^{p}_{+}=\partial\mathbb{V}^{p}_{+}$ is the infinitesimally extended surface domain of a particle $p$ and $\bm{n}^{p}_{+}$ its outward-directed unit normal vector. The bottom line of Eq. (45) is precisely how $\bm{F}^{p}_{\mathrm{h}}$ is calculated in Uhlmann’s IBM [17, 18, 19, 20, 21], considering that these studies modeled the surface $\mathbb{S}^{p}$ as a very thin volume layer, the so-called Lagrangian layer [20], extending minimally both into the interior and exterior of the particle $p$. This layer was then discretized into small grid cells, so that in total:
@@
where $\bm{f}^{p}_{l}\Delta V^{p}_{l}$ is the force applied on the grid cell volume $\Delta V^{p}_{l}$ corresponding to a given Lagrange point (index $l$) on the surface of the particle $p$. In Eq. (46), and also in other equations, such as those of Appendix C, the symbol $\simeq$’ shall indicate approximations associated with DNS- or IBM-related discretizations.
@@
Here $\mathcal{G}_{\Delta}[\bm{x}]$ is a smoothing function that satisfies $\int_{\mathbb{R}^{3}}\mathcal{G}_{\Delta}[\bm{x}]\mathrm{d}x^{3}=1$ and exhibits a compact support on the order of the DNS grid length $\Delta$ (Appendix B). The key problem now is that $\mathcal{G}_{\Delta}$ spreads the particle surface force density $\bm{\tilde{f}}^{p}_{\mathrm{IBM}}$ onto the stationary Eulerian DNS grid, independently of $\Delta$ about equally to the interior and exterior of $\mathbb{V}^{p}_{+}$, with the result that $|\int_{\mathbb{V}^{p}_{+}}\overline{\bm{f}}_{\mathrm{IBM}}\mathrm{d}V|$ substantially underestimates $|\int_{\mathbb{V}^{p}_{+}}\bm{f}_{\mathrm{IBM}}\mathrm{d}V|$, leading to potentially substantial differences between $\int_{\mathbb{S}^{p}_{+}}\bm{n}^{p}_{+}\cdot\bm{\tilde{\sigma}}\mathrm{d}S$ and $\bm{F}^{p}_{\mathrm{h}}$. This issue remains present even in the limit $\Delta\to 0^{+}$, implying that it cannot be dismissed as a discretization error. It suggests that Eq. (47) may not be well approximated when the unknown $\bm{\sigma}$ on its right-hand side is approximated by $\bm{\tilde{\sigma}}$. In addition, in order to evaluate Eq. (47), one would then need to interpolate the values of $\bm{\tilde{\sigma}}$ back from the DNS grid to the particle surfaces, creating another potential source of inaccuracy.
@@
To solve both problems, we exploit the fact that the fluid fields are well-defined in the entire fluid-particle domain $\mathbb{V}_{\infty}$. Using the relations derived in Appendix A and $\bm{\nabla}\cdot\bm{\sigma}=\bm{\nabla}\cdot\bm{\tilde{\sigma}}+\overline{\bm{f}}_{\mathrm{IBM}}-\bm{f}_{\mathrm{IBM}}$, this permits calculating [^2] The last equality in Eq. (50) neglects the fact that particles are allowed to very slightly overlap. It is possible to account for particle overlap through replacing $S^{p}\underline{\bm{r}^{p}\bm{\tilde{f}}^{p}_{\mathrm{IBM}}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{S}^{p}}$ by $X^{p}_{+}\sum\nolimits_{q}S^{q}\underline{\bm{r}^{p}\bm{\tilde{f}}^{q}_{\mathrm{IBM}}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{S}^{q}}$ in Eq. (50), and subsequently in Eq. (53), which gives nonzero contributions not only for $q=p$ but also for $q\neq p$ if a portion of the surface $\mathbb{S}^{q}$ of a particle $q$ is within the extended domain $\mathbb{V}^{p}_{+}$ of a particle $p$. However, this would be inconsistent with the fact the we already neglected particle overlap in Eq. (45). We confirmed that, for our simulations presented in §3.2, the effect of particle overlap on $\bm{\sigma^{\mathrm{f}}}$ is indeed negligible.
@@
$\begin{split}&\bm{\nabla}\cdot S^{p}\underline{\bm{r}^{p}\bm{n}^{p}_{+}\cdot\bm{\sigma}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{S}^{p}_{+}}\\ =&\;\bm{\nabla}\cdot\left(X^{p}_{+}\bm{\sigma}+V^{p}\underline{\bm{r}^{p}(\bm{\nabla}\cdot\bm{\sigma})\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{V}^{p}_{+}}\right)\\ =&\;\bm{\nabla}\cdot\left(X^{p}_{+}\bm{\sigma}+V^{p}\underline{\bm{r}^{p}\left(\bm{\nabla}\cdot\bm{\tilde{\sigma}}+\overline{\bm{f}}_{\mathrm{IBM}}\right)\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{V}^{p}}\right.\\ &\quad\quad\quad\quad\;\;\,-\left.S^{p}\underline{\bm{r}^{p}\bm{\tilde{f}}^{p}_{\mathrm{IBM}}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{S}^{p}}\right),\end{split}$ (50)
@@
Furthermore, since $\mathcal{G}_{\Delta}$ integrates to unity, the difference $\bm{\nabla}\cdot\bm{\sigma}-\bm{\nabla}\cdot\bm{\tilde{\sigma}}$ can be directly calculated from the surface force density $\bm{\tilde{f}}^{p}_{\mathrm{IBM}}$ as
@@
$\begin{split}&\bm{\sigma^{\mathrm{f}}}=\langle\bm{\tilde{\sigma}}\rangle+\bm{\sigma^{\mathrm{f}}_{\mathrm{Re}}}+\bm{\sigma^{\mathrm{f}}_{\mathrm{\tilde{s}}}}+\bm{\sigma^{\mathrm{f}}_{\Delta}}\\ &~~~\,\simeq\langle\bm{\tilde{\sigma}}\rangle+\bm{\sigma^{\mathrm{f}}_{\mathrm{Re}}}+\bm{\sigma^{\mathrm{f}}_{\mathrm{\tilde{s}}}},\end{split}$ (52)
$\begin{split}&\bm{\sigma^{\mathrm{f}}_{\mathrm{\tilde{s}}}}\equiv\left\langle\sum\nolimits_{p}\left(V^{p}\underline{\bm{r}^{p}\left(\bm{\nabla}\cdot\bm{\tilde{\sigma}}+\overline{\bm{f}}_{\mathrm{IBM}}\right)\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{V}^{p}}-\right.\right.\\ &\left.\left.S^{p}\underline{\bm{r}^{p}\bm{\tilde{f}}^{p}_{\mathrm{IBM}}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{S}^{p}}-\underline{\bm{r}^{p}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathbb{V}^{p}}\bm{F}^{p}_{\mathrm{h}}\right)\right\rangle,\end{split}$ (53)
@@
$\begin{split}\bm{\sigma^{\mathrm{m}}}&=\langle\bm{\tilde{\sigma}}\rangle+\bm{\sigma^{\mathrm{f}}_{\mathrm{\tilde{s}}}}+\bm{\sigma^{\mathrm{f}}_{\Delta}}+\bm{\sigma^{\mathrm{s}}_{\mathrm{c}}}+\bm{\sigma^{\mathrm{m}}_{\mathrm{Re}}}\\ &\simeq\langle\bm{\tilde{\sigma}}\rangle+\bm{\sigma^{\mathrm{f}}_{\mathrm{\tilde{s}}}}+\bm{\sigma^{\mathrm{s}}_{\mathrm{c}}}+\bm{\sigma^{\mathrm{m}}_{\mathrm{Re}}}.\end{split}$ (55)
In Eqs. (52) and (55), we neglected the contribution $\bm{\sigma^{\mathrm{f}}_{\Delta}}$, since it vanishes in the limit $\Delta\to 0^{+}$ due to $\mathcal{G}_{\Delta}[\bm{r}]\to\delta[\bm{r}]$, meaning that $\bm{\sigma^{\mathrm{f}}_{\Delta}}$ is a true artifact of the DNS- and IBM-related discretizations [^3] We confirmed that, for our simulations presented in §3.2, $\bm{\sigma^{\mathrm{f}}_{\Delta}}$ is indeed negligible.. Hence, the crucial effect of the IBM-related smoothing with the function $\mathcal{G}_{\Delta}$ in Eqs. (48) and (49) is the intermingling of the Eulerian volumetric force density field $\overline{\bm{f}}_{\mathrm{IBM}}$ and the singular surface force $\bm{\tilde{f}}^{p}_{\mathrm{IBM}}$ per unit area in Eq. (53), since this effect does not vanish in the limit $\Delta\to 0^{+}$.